Demand is P=70-0.5Q, MC=10.

(a) What is optimal price quantity if firm charges only one price?

(b) What is the price quantity and extra profit firm can earn if firm use blocking price where the first block is sold at price you determined in the previous question. ***I believe that this part is referring to the price in part a***

(c) Can firm be more profitable using a two block price other than what you determined here?
Answer #1

a) Since the demand is downward sloping, the firm has market power and so the optimal price quantity equilibrium is determined at MR = MC. Here MR is 70 - Q so we have 70 - Q = 10 and Q = 60 units. This gives price P = 70 - 60*0.5 = $40. Thus, when only one price is charged, the price is $40 and quantity is 60 units

b) The block size is now 60 units. Firm can charge a block price which is equivalent to the sum of consumer surplus and sales revenue at this unit. The block price is 0.5*(70 - 40)*60 + 40*60 = 3300. Quantity is 60 units. Extra profit = New profit - old profit = (3300 - 10*60) - (40*60 - 10*60) = 900

c) Yes. The firm can use many blocks and as the number of blocks increases, the profit will increase.